OpenStudy (rose4825):

How many vertices does a convex polyhedron have if it has 14 faces and 24 edges? 10 12 11 8 My answer: 12

OpenStudy (rose4825):

@mathmale

OpenStudy (mathstudent55):

Isn't there a formula for this?

OpenStudy (rose4825):

I believe it is V+F=E+2

OpenStudy (mathstudent55):

https://plus.maths.org/content/eulers-polyhedron-formula

OpenStudy (rose4825):

I just want to double check my answers

OpenStudy (mathstudent55):

Correct.

OpenStudy (rose4825):

Can you check a few more?

jimthompson5910 (jim_thompson5910):

either that or V-E+F = 2 but that's equivalent to what you wrote @rose4825 https://plus.maths.org/content/eulers-polyhedron-formula

OpenStudy (rose4825):

What is the Euler characteristic of a polyhedron that has 12 vertices, 20 faces, and 30 edges? 22 12 8 2 My answer: 2

OpenStudy (mathstudent55):

correct

OpenStudy (rose4825):

I will open a new question :) so I can give other medals